# Addressing Sample Inefficiency in Multi-View Representation Learning

### v1

@@ Abstract
Non-contrastive self-supervised learning (NC-SSL) methods like BarlowTwins and VICReg have shown great promise for label-free representation learning in computer vision. Despite the apparent simplicity of these techniques, researchers must rely on several empirical heuristics to achieve competitive performance, most notably using high-dimensional projector heads and two augmentations of the same image. In this work, we provide theoretical insights on the implicit bias of the BarlowTwins and VICReg loss that can explain these heuristics and guide the development of more principled recommendations. Our first insight is that the orthogonality of the features is more critical than projector dimensionality for learning good representations. Based on this, we empirically demonstrate that low-dimensional projector heads are sufficient with appropriate regularization, contrary to the existing heuristic. Our second theoretical insight suggests that using multiple data augmentations better represents the desiderata of the SSL objective. Based on this, we demonstrate that leveraging more augmentations per sample improves representation quality and trainability. In particular, it improves optimization convergence, leading to better features emerging earlier in the training. Remarkably, we demonstrate that we can reduce the pretraining dataset size by up to 4x while maintaining accuracy and improving convergence simply by using more data augmentations. Combining these insights, we present practical pretraining recommendations that improve wall-clock time by 2x and improve performance on CIFAR-10/STL-10 datasets using a ResNet-50 backbone. Thus, this work provides a theoretical insight into NC-SSL and produces practical recommendations for enhancing its sample and compute efficiency.
@@ 1 Introduction
Unsupervised representation learning, i.e., learning features without human-annotated labels, is critical for progress in computer vision. Modern approaches, grouped under the self-supervised learning (SSL) umbrella, build on the core insight that similar images should map to nearby points in the learned feature space. Current SSL methods can be broadly categorized into contrastive and non-contrastive algorithms. While both categories aim to learn the desired features using “positive” samples, which refer to different augmentations of the same image, they diverge in using “negative” samples. Contrastive methods use augmentations obtained from entirely different images as negative samples to avoid the trivial solution of mapping all samples to the same point in the feature space (i.e., representational collapse). But, this necessitates an elaborate sampling scheme and huge batch sizes. Non-contrastive methods, on the other hand, eliminate the need for negative samples altogether and instead rely on regularizing the feature space to avoid representational collapse.
A prominent subgroup among non-contrastive SSL methods is the family of Canonical Correlation Analysis (CCA) algorithms, which includes BarlowTwins [32] and VICReg [5]. These methods aim to enforce orthogonality among the learned features in addition to learning to map similar images to nearby points in feature space and have been shown to achieve competitive performance on benchmark computer vision datasets. These methods have become the preferred strategy for representation learning in several domains due to the lack of need for negative samples and their simple formulation. However, despite the apparent simplicity of their loss functions, the behavior of this family of algorithms is not well understood. Therefore, researchers often use empirically driven heuristics to design successful applications, such as using (i) a high-dimensional projector head or (ii) two augmentations per image. Although these heuristics help in practice, their theoretical underpinnings are unclear.
Figure 1: Design of existing SSL algorithms relies on heuristics. (A) We investigate the theoretical underpinnings of (i) the number of augmentations and (ii) the dimensionality of the projector $g_{\phi}$. (B) We show that a generalized NC-SSL loss with multiple augmentations, & low-dimensional projectors outperforms existing heuristics while using 4$\sim$ 4$\times$ 4 × smaller pretraining dataset.
Alongside relying on heuristics and researchers’ intuition for design, existing SSL algorithms are extremely data-hungry. In particular, state-of-the-art algorithms often rely on large-scale datasets [26] or data engines [23] to achieve good representations. While this strategy works exceptionally well in natural-image settings, its application is limited in other critical domains, such as medical imaging, where the number of samples is scarce.
With these challenges in mind, the primary focus of this work is making progress toward establishing theoretical foundations underlying the family of non-contrastive SSL algorithms (NC-SSL) with an eye toward sample efficiency. In particular, we analyze the BarlowTwins and VICReg losses and show that they implicitly learn the data similarity kernel defined by the chosen augmentations. We show that learning the data similarity kernel is helped by greater orthogonality in the projector outputs and more data augmentations. As such, increasing the orthogonality of the projector output eliminates the requirement for a high-dimensional projector head, and increasing the number of data augmentations decreases the number of unique samples required. Our theoretical analysis establishes a principled grounding for the role of multiple augmentations and the sufficiency of low-dimensional projectors, outlining a framework for improving the sample efficiency of NC-SSL while maintaining representation quality.
@@
We empirically verify our theoretical insights using the popular ResNet-50 backbone on benchmark datasets, CIFAR-10 and STL-10. Strikingly, we show that our multi-augmentation approach can learn good features even with a quarter of the samples in the pretraining dataset. As such, this suggests that SSL training can be done with smaller datasets and opens exciting questions in the design of performance-enhancing transformations. In summary, our core contributions are:
  • Eigenfunction interpretation: We demonstrate that the loss functions of the CCA family of non-contrastive SSL algorithms are equivalent to the objective of learning eigenfunctions of the augmentation-defined data kernel.
@@
  • Role of heuristics: We provide a mechanistic explanation for the role of projector dimensionality and the number of data augmentations and empirically demonstrate that low-dimensional projector heads are sufficient and that using more augmentations leads to learning better representations.
@@
  • Data efficient NC-SSL: Leveraging the convergence benefits of the multi-augmentation framework, we demonstrate that we can learn good features with significantly smaller datasets (up to 25%) without harming downstream performance.
2 Related Work
Self-Supervised Pretraining requires significant compute resources, and with the lack of a unified theoretical framework, most practitioners rely on empirical heuristics. The SSL cookbook [3] provides a comprehensive summary of several widely adopted practices. While recent advances in SSL theory explore learning dynamics in linear (or shallow) models [30, 29], with a focus on understanding dimensionality collapse [20, 15], the theoretical underpinnings of most “recipes” essential for good feature learning, are missing.
Contrastive SSL has received more theoretical attention, owing to its connection with metric learning and noise contrastive estimation [22, 4, 19]. In particular, [18] provides a theoretical framework for the SimCLR loss from an augmentation graph perspective, which leads to practical recommendations. Subsequently, [14] establish a duality between contrastive and non-contrastive learning objectives, further bridging the gap between theory and practice.
Non-contrastive SSL algorithms have comparatively scarce theoretical foundations. In prior work [1, 14] find that with modified learning objectives, low-dimensional projectors are sufficient for good downstream performance. Similarly, previous works have demonstrated notable performance boosts when using a multi-patch framework in contrastive [10] and non-contrastive SSL. However, the theoretical basis for the benefits and trade-offs of either low-dimensional projectors or multiple augmentations is unclear.
Deep Learning theory has made significant strides in understanding the optimization landscape and dynamics of supervised learning [2]. In concurrent works, [33, 7] explore the interplay between the inductive bias of data augmentations, architectures, and generalization. In recent work, [28] uses a more straightforward formulation of the BarlowTwins loss and investigates the learning dynamics in linearized models for the case when the invariance and orthogonalization losses have equal penalties.
@@ 3 Preliminaries
We start by formally defining the unsupervised representation learning problem for computer vision. In particular, we assume access to a dataset $\mathcal{D}=\{{x_{1},x_{2},...,x_{n}}\}$ with $x_{i}\in\mathbb{R}^{p}$ consisting of unlabeled instances (often natural images), and the objective is to learn a $d$-dimensional representation ($d<p$) that is useful across multiple downstream applications. We focus on learning the parameters of a deep neural network $f_{\theta}\in\mathcal{F}_{\Theta}$, using the multi-view invariance SSL framework, wherein multiple views of an example are used to optimize pretraining loss function, $\mathcal{L}_{pretrain}(f_{\theta},\mathcal{D})$
Non-Contrastive Self-Supervised Learning (NC-SSL) algorithms impose invariance to data-augmentations, which are used to define multiple views of the same image while imposing specific regularization on the geometry of the learned feature space. More generally, both families can be thought of as decomposing $\mathcal{L}_{pretrain}$ into two terms (i) $\mathcal{L}_{invariance}$: to learn invariance to data augmentations and (ii) $\mathcal{L}_{collapse}$ to prevent collapsing the feature space to some trivial solution with no discriminative power.
@@
where $\beta$ denotes a hyperparameter that controls the importance of the collapse-preventing term relative to the invariance term.
This formulation separates instance-level attributes invariant to augmentations, highlighting the semantic information of the instance. The ideal feature space is less sensitive to varying attributes and more sensitive to semantic ones, facilitating generalization to new examples. Understanding the interplay between pretraining loss and preserved attributes is critical for time and compute efficiency.
Data Augmentation graph was introduced by [18] to analyze contrastive losses, like SimCLR [8]. Briefly, we define a graph $\mathcal{G(X,W)}$ with the vertex set ($\mathcal{X},\rho_{X}$) comprising the result of all possible data augmentations from each sample in a dataset (could be infinite when continuous augmentation functions are used) and $\mathcal{W}$ denoting the adjacency matrix. Let $x_{0}$ be an image in $\mathcal{X}$, and let $z=M(x_{0})$ be a random data augmentation of the image, $x_{0}$. We define the probability density of reaching $z$ from $x_{0}$ via a choice of mapping $M$:
@@
Since mapping is not generally invertible (e.g., crops), observe that $p(x_{0}\mid z)\not=p(z\mid x_{0})$. Using this definition, we now define the strength of the edge between nodes $x$ and $z$ of the augmentation graph as the joint probability of generating augmentations $x,z$ from the same image $x_{0}\sim\rho_{X}$. Formally,
$w_{xz}:=\mathbb{E}_{x_{0}\sim\rho_{X}}\left[p(x\mid x_{0})p(z\mid x_{0})\right]$ (3)
It is worth noting that the magnitude of $w_{xz}$ captures the relative similarity between $x$ and $z$. A higher value of $w_{xz}$ indicates that both patches are more likely to come from the same image and, thereby, more similar.
Figure 2: Schematic of augmentation graph. (A) Augmentations from each image span a region in the image space, which could overlap with the augmentation span of other images. (B) An augmentation graph schematic that uses probabilities to characterize the interactions among augmentation spans of different instances.
[18] showed that optimizing a functionally equivalent form of the SimCLR loss termed the spectral contrastive loss ($\mathcal{L}_{c}$, essentially learns features whose covariance structure matches the adjacency matrix of the augmentation graph.
$\mathcal{L}_{c}\propto\|ZZ^{T}-\bar{\mathcal{W}}\|^{2}_{F}$ (4)
where $Z$ denotes the output of the neural network, $\bar{\mathcal{W}}$ denotes the degree-normalized adjacency matrix and .$\|.\|_{F}$ denotes the Frobenius norm operator. This perspective implies that the features learned by a contrastive SSL framework would align with the top eigenvectors of $\bar{\mathcal{W}}$. As observed by [18], all rotations of $Z$ that don’t change its span define an equivalence class of solutions to the above optimization problem and make no difference for the downstream generalization of a linear probe. Based on this insight, we define a notion of equivalence among learned feature spaces.
@@
4 Data augmentation kernel perspective of non-contrastive SSL
Following the previous section, we present an augmentation kernel perspective of BarlowTwins and VICReg losses. Specifically, we show that these losses are equivalent to the optimization problem of learning eigenfunctions of the augmentation-defined data covariance kernel. Subsequently, we argue that using a high-dimensional projector yields better overlap with the top eigenvectors of the data augmentation kernel at initialization compared to a low-dimensional projector. Therefore, our analysis suggests using a stronger orthogonalization constraint during optimization for lower-dimensional projectors to ensure that features learned are equivalent to those learned with high-dimensional projectors. Furthermore, we argue that using more augmentations improves our estimate of the augmentation-defined data covariance kernel, thereby aiding the eigenfunction optimization problem. Therefore, our analysis suggests using an averaging operator with more data augmentations to estimate the true augmentation kernel better.
@@ 4.1 Features in terms of data augmentation kernels
We will define two notions of the data augmentation kernel. Given two images, $x,z$, the first kernel, which we call the forward data augmentation covariance kernel, is given by
$k^{DAF}(x,z)=\mathbb{E}_{x_{0}\sim\rho_{X}}\left[\frac{p(x\mid x_{0})}{\rho(x)}\frac{p(z\mid x_{0})}{\rho(z)}\right]$ (6)
This covariance kernel measures the similarity between $x,z$ in terms of how likely they are to be reached from $x_{0}$, weighted by the distribution of $x_{0}$. Note that this is the edge strength between nodes $x,z$ in the augmentation graph. We can also define a (backward) data augmentation covariance kernel that reverses the roles of ($x$,$z$) and $x_{0}$:
$k^{DAB}(x,z)=\mathbb{E}_{x_{0}\sim\rho_{X}}\left[\frac{p(x_{0}\mid x)}{\rho(x_{0})}\frac{p(x_{0}\mid z)}{\rho(x_{0})}\right]$ (7)
SSL aims to learn features that preserve the covariance kernel structure (imposed by this choice of mapping $M$) [11]. Therefore, we want to define a loss which determines vector features, $F:X\to\mathbb{R}^{d}$, which factor a data augmentation kernel $k^{DA}(x,z)=F(x)^{\top}F(z)$. Doing this directly is prohibitively data-intensive at scale since it involves a search over data-augmented images. However, since the covariance kernels are PSD, they define a Reproducing Kernel Hilbert space (RKHS). This allows us to apply Mercer’s theorem to find vector features as in [12, 13, 24].
The construction of features using Mercer’s theorem goes as follows. Given a PSD data augmentation kernel, $k^{DA}$, define the $T_{k}$ operator, which takes a function $f$ and returns its convolution with the data augmentation kernel.
@@
We will also make use of the following operator,
@@
which averages the values of the function, $f$, over the augmented images $x_{0}=M(x)$ of the data, $x$.
Since the operator $T_{k}$ is compact and positive, it has a spectral decomposition consisting of eigenfunctions $\phi_{i}$ and corresponding eigenvalues $\lambda_{i}$. Using these eigenpairs, we can define the (infinite sequence of square summable) spectral features, $G:X\to\ell_{2}$, (where $\ell_{2}$ represents square summable sequences), by
$G(x)=(\sqrt{\lambda_{1}}\phi_{1}(x),\dots,\sqrt{\lambda_{d}}\phi_{d}(x),\dots)$ (10)
Then, Mercer’s theorem gives
$k^{DA}(x,z)=G(x)\cdot G(z)$ (Mercer)
and ensures that the inner product is finite. These are the desired features that factor in the kernel. However, computing the eigenfunctions of $T_{k}$ is costly. Instead, we propose an alternative using the more efficient operator $T_{M}$. According to Definition 3.1, both operators lead to equivalent features.
@@
The idea of the proof uses the fact that, as linear operators, $T_{k^{DAB}}=T_{M}^{\top}T_{M}$ and that $T_{k^{DAF}}=T_{M}T_{M}^{\top}$. Then we use the spectral theory of compact operators, which is analog of the Singular Value Decomposition in Hilbert Space, to show that eigenfunctions of $T_{M}^{\top}T_{M}$ operator are the same as those obtained from optimizing $L(F)$. A similar result can be obtained using $k^{DAF}$ and $T_{M}^{\top}$.
4.2 Corollary 1: Low-dimensional projectors are sufficient

### v2

# Harnessing small projectors and multiple views for efficient vision pretraining

@@ Abstract
Recent progress in self-supervised (SSL) visual representation learning has led to the development of several different proposed frameworks that rely on augmentations of images but use different loss functions. However, there are few theoretically grounded principles to guide practice, so practical implementation of each SSL framework requires several heuristics to achieve competitive performance. In this work, we build on recent analytical results to design practical recommendations for competitive and efficient SSL that are grounded in theory. Specifically, recent theory tells us that existing SSL frameworks are actually minimizing the same idealized loss, which is to learn features that best match the data similarity kernel defined by the augmentations used. We show how this idealized loss can be reformulated to a functionally equivalent loss that is more efficient to compute. We study the implicit bias of using gradient descent to minimize our reformulated loss function, and find that using a stronger orthogonalization constraint with a reduced projector dimensionality should yield good representations. Furthermore, the theory tells us that approximating the reformulated loss should be improved by increasing the number of augmentations, and as such using multiple augmentations should lead to improved convergence. We empirically verify our findings on CIFAR, STL and Imagenet datasets, wherein we demonstrate an improved linear readout performance when training a ResNet-backbone using our theoretically grounded recommendations. Remarkably, we also demonstrate that by leveraging these insights, we can reduce the pretraining dataset size by up to 2$\times$ while maintaining downstream accuracy simply by using more data augmentations. Taken together, our work provides theoretically grounded recommendations that can be used to improve SSL convergence and efficiency.
@@ 1 Introduction
Unsupervised representation learning, i.e., learning features without human-annotated labels, is critical for progress in computer vision. Modern approaches, grouped under the self-supervised learning (SSL) umbrella, build on the core insight that similar images should map to nearby points in the learned feature space – often termed as the invariance criterion. Current SSL methods can be broadly categorized into contrastive and non-contrastive algorithms, based on whether they formulate their loss functions using negative samples or not, respectively.
Despite this difference in their loss formulations, recent theoretical work has established an equivalence between the contrastive and non-contrastive SSL frameworks [19]. This work shows that these different SSL formulations are ultimately minimizing a loss that encourages the learning of features that best match the data similarity kernel defined by the augmentations used. However, this notion of theoretical equivalence holds only in the limit of ideal pretraining settings, i.e. with access to infinite data and compute budget, and the feature learning behavior of different SSL algorithms in practical scenarios is still not well understood. Therefore, researchers often use empirically driven heuristics that are theoretically ungrounded to design successful applications, such as (i) a high-dimensional projector head for non-contrastive SSL or (ii) the use of two augmentations per image [3]. Moreover, existing SSL algorithms are extremely data-hungry, relying on large-scale datasets [33] or data engines [30] to achieve good representations. While this strategy works exceptionally well in data-rich settings (like training on natural-images), it is not viable in data-constrained settings (like medical imaging), where samples are relatively scarce.
Figure 1: Design of existing SSL algorithms relies on heuristics. (A) Augmentation graphs are common in vision pretraining, providing generalizable features for downstream tasks. (B) We propose an equivalent loss function for SSL pretraining that recovers the same eigenfunctions more efficiently than existing approaches.
With these challenges in mind, the primary focus of this work is to develop theoretically grounded recommendations for improving the effectiveness and efficiency of feature learning, both with respect to the required compute budget as well as data points. Like any unsupervised representation learning algorithm, features learned through SSL depend on three factors: (i) implicit bias of the architecture, (ii) explicit invariance imposed by data augmentations, (iii) implicit bias of the learning rule. While previous works predominantly studied the role of the model architecture capacity and loss function, and their interplay with data augmentations [9, 44], our approach broadens this perspective by also considering the role of the learning rule (gradient descent) in optimizing these loss functions. Specifically, we extend the previous theoretical findings [44] that unified the desiderata of different SSL algorithms. We reformulate the idealized unifying loss to propose a functionally equivalent loss that is more compute-efficient (see Figure 1). Based on our loss formulation, we provide two practical recommendations that can help improve the efficiency of SSL pipelines while maintaining good performance. First, we show that optimizing the reformulated loss using gradient descent can often reduce the orthogonality among the learned embeddings, thereby leading to an inefficient use of the projector network’s capacity. Consequently, we recommend using a stronger orthogonalization constraint to eliminate the requirement of high-dimensional projector heads, thereby significantly reducing the parameter overhead of good feature learning. Second, we show that increasing the number of augmentations leads to a better estimate of the data similarity kernel. Consequently, we recommend using more augmentations to improve optimization convergence and learn better features earlier in training.
@@
We empirically verify our theoretically grounded recommendations using the popular ResNet backbone on benchmark datasets: CIFAR, STL and Imagenet. Strikingly, we show that our multi-augmentation approach can learn good features even with half of the samples in the pretraining dataset. Our recommendations provide a path towards making SSL pretraining more data and compute-efficient without harming performance and could unlock massive performance gains in data-constrained setups. In summary, our core contributions are as follows:
  • Efficient SSL loss formulation: We propose an functionally equivalent and compute-efficient formulation of the SSL desiderata that yields the eigenfunctions of the augmentation-defined data similarity kernel.
@@
  • Role of heuristics: Based on our loss formulation and the implicit bias of gradient descent in optimizing this loss, we provide a mechanistic explanation for the role of projector dimensionality and the number of data augmentations. Consequently, we empirically demonstrate that low-dimensional projector heads are sufficient and that using more augmentations leads to learning better representations.
@@
  • Data efficient SSL: Leveraging the convergence benefits of the multi-augmentation SSL framework, we empirically demonstrate that we can learn good features with significantly smaller datasets (up to 2×2$\times$2 ×) without harming downstream linear probe performance.
@@ 2 Preliminaries
Existing SSL approaches in computer vision In recent years machine learning researchers have developed a number of effective approaches for learning from data without labels. The most popular approaches use augmentations of data points as targets for themselves. One of the first was a Simple framework for Contrastive Learning (SimCLR), which relied on an infoNCE loss with augmentations of an image as positive targets and augmentations of other images as negative samples (to construct the contrastive loss) [12]. Other works have relied on non-contrastive approaches, notably BarlowTwins [43] and VICReg [5]. BarlowTwins, which was inspired by the ideas of the neuroscientist Horace Barlow (cite), also uses augmentations of images, but it instead aims to optimize the covariance structure of the representations in order to reduce redundancies in the feature space [43]. Variance Invariance Covariance Regularization (VICReg) was a modification of BarlowTwins that added a variance term in the loss in order to ensure that every feature dimension has a finite variance [5]. In this paper we will focus on non-contrastive methods like BarlowTwins and VICReg, but in line with previous work [44], we also consider how these approaches relate to contrastive methods like SimCLR.
Formalizing the self-supervised learning problem Now, we will formalize the unsupervised representation learning problem for computer vision. In particular, we assume access to a dataset $\mathcal{D}=\{{x_{1},x_{2},...,x_{n}}\}$ with $x_{i}\in\mathbb{R}^{p}$ consisting of unlabeled images. The objective is to learn a $d$-dimensional representation ($d<p$) that is useful across multiple downstream applications. We focus on learning the parameters of a deep neural network $f_{\theta}\in\mathcal{F}_{\Theta}$, using the multi-augmentation SSL framework, wherein multiple views of an image are used to optimize the pretraining loss function, $\mathcal{L}_{pretrain}(f_{\theta},\mathcal{D})$
Non-Contrastive Self-Supervised Learning (NC-SSL) algorithms impose invariance to data augmentations, while imposing regularization on the geometry of the learned feature space. More generally, $\mathcal{L}_{pretrain}$ can be formulated with two terms (i) $\mathcal{L}_{invariance}$: to learn invariance to data augmentations and (ii) $\mathcal{L}_{collapse}$: regularization to prevent collapsing the feature space to a trivial solution.
@@
where $\beta$ denotes a hyperparameter that controls the importance of the collapse-preventing term relative to the invariance term. This formulation separates features that are invariant to the augmentations from those that are sensitive to them. Intuitively, the ideal feature space is more sensitive to semantic attributes (e.g. “that’s a dog”) and less sensitive to irrelevant attributes (e.g. “direction the dog is facing”), facilitating generalization to new examples.
Data Augmentation graph was introduced by Haochen et al. [22] to analyze contrastive losses, like SimCLR [12]. Briefly, we define a graph $\mathcal{G(A,W)}$ that captures the relationship between images derived from all possible data augmentations. The vertex set ($\mathcal{A},\rho_{A}$) is each augmented sample in a dataset, $\mathcal{X}$, and the adjacency matrix, $\mathcal{W}$, denotes the similarity between pairs of vertices. Let $x_{0}$ be an image in $\mathcal{X}$, and let $z=M(x_{0})\in\mathcal{A}$ be a random data augmentation of the image, $x_{0}$. We define the probability density of reaching $z$ from $x_{0}$ via a choice of mapping $M$:
@@
Since the mapping is not generally invertible (e.g., cropping), we observe that $p(x_{0}\mid z)\not=p(z\mid x_{0})$. Using this definition, we now formally define the strength of the edge between nodes $x,z\in\mathcal{A}$ of the augmentation graph as the joint probability of generating augmentations $x,z$ from the same image $x_{0}\sim\rho_{X}$. Notably, the edge strength of the (degree-normalized) augmentation graph is equivalent to the data similarity kernel, defined in [44]. Formally,
$k^{DAF}(x,z)=w_{xz}:=\mathbb{E}_{x_{0}\sim\rho_{X}}\left[\frac{p(x\mid x_{0})}{p(x)}\frac{p(z\mid x_{0})}{p(z)}\right]$ (3)
The magnitude of $w_{xz}$ captures the augmentation-defined similarity between $x$ and $z$. A higher value of $w_{xz}$ indicates that both patches are more likely to come from the same image and, thereby, are more similar.
The desiderata of different SSL algorithms can be understood as learning features $F$ that best capture $k^{DAF}(x,z)$, i.e. $F(x)^{T}F(z)\approx k^{DAF}(x,z)$. Recent theoretical work has shows that different SSL losses can be formulated as special cases of the objective function that recovers the top-d eigenfunctions of $k^{DAF}(x,z)$ [44].
$\mathcal{L}_{ssl}(F)=\mathbf{E}_{x,z\in\mathcal{A}}\left[(k^{DAF}(x,z)-F(x)^{T}F(z))^{2}\right]$ (4)
Note that all rotations of $F$ that don’t change its span define an equivalence class of solutions to Equation 4 and make no difference for the downstream generalization of a linear probe. Based on this insight, we define an equivalence among learned feature spaces:
@@
3 Implicit bias of non-contrastive SSL loss and optimization
We extend the recent theoretical results [44] to propose a compute-efficient reformulation of the loss function of the SSL desiderata that yields equivalent features, i.e. the functions spanning the eigenfunctions of the augmentation-defined data similarity kernel, $k^{DAF}$. Furthermore, we study the role of gradient descent in optimizing this loss function and uncover a selection and primacy bias in feature learning. Specifically, we find that gradient descent tends to learn the dominant eigenfunctions (eigenfunctions corresponding to larger eigenvalues) earlier during training, and often over-represents these eigenfunctions under weak orthogonalization constraints.
Consequently, we propose employing a stronger orthogonalization constraint during optimization when using a low-dimensional projector to ensure that learned features are equivalent to those learned with a high-dimensional projector. Furthermore, we argue that using more augmentations improves our sample estimate of $k^{DAF}$, thereby aiding the eigenfunction optimization problem. We dedicate the rest of this section to highlight our key theoretical insights, and practical recommendations that follow them.
@@ 3.1 Features in terms of data augmentation kernels
Let us define a kernel operator, $T_{k}$, for a positive semi-definite data augmentation kernel, $k^{DAF}$.
@@
such that Equation 4 can be equivalently written as (Equation 5 of [44])
$\mathcal{L}_{ssl}(F)=\langle F,(I-T_{k})F\rangle_{\rho_{\mathcal{A}}}$ (7)
We can now use Mercer’s theorem to factorize $k^{DAF}$ into corresponding spectral features $G:X\to\ell_{2}$ (where $\ell_{2}$ represents square summable sequences) [15, 16, 31]. However, note that computing $k^{DAF}$ (or $T_{k}$) is expensive as it requires computing the overlap among all augmentations of every pair of data points. Instead of computing the eigenfunctions of $T_{k}$ directly, we propose using an alternative operator $T_{M}$:
@@
which averages the values of the function, $f$, over the augmented images $x_{0}=M(x)$ of the data, $x$. We show that $T_{M}^{T}T_{M}$ is equivalent to $T_{k}$, and therefore $T_{M}$ and $T_{k}$ have shared eigenfunctions.
@@
Let $G(x)$ be the infinite Mercer features of the backward data augmentation covariance kernels, $k^{DAB}$. Let $F(x)=(f_{1}(x),\dots,f_{N_{k}}(x))$ be the features given by minimizing the following data augmentation invariance loss
@@
which includes the orthogonality constraint. Then, $V(F)\subset V(G)$ , $\lim_{N_{k}\to\infty}V(F)=V(G)$.
As shown in the Appendix B, $L(F)$ is equivalent to a constrained optimization formulation of the BarlowTwins loss. Furthermore, $L(F)$ with the additional constraint that $(f_{i},f_{i})\geq\gamma$ $\forall i\in\{1,2\dots N_{k}\}$ is the constrained optimization formulation of the VICReg loss.
3.2 The implicit bias of gradient descent
Next, we investigate how the use of gradient descent for optimizing $L(F)$ influences the characteristics of the learned feature space, $V(F)$. Given the similarity in its form with that of the BarlowTwins loss, we build on recent findings that demonstrate the sequential nature of learning eigenfunctions when optimizing the BarlowTwins loss under a strong orthogonalization regularization [36]. Since strong orthogonalization is seldom used in practice due to instabilities in training [43, 5], we believe studying the learning dynamics under weak orthogonalization regularization (i.e. low values of $\beta$ in Equation 1) is more relevant to provide recommendations for practitioners.
Theorem 3.2.
(Informal) Let us denote the span of the feature space at initialization as $V(F_{0})$ and after training as $V(F_{T})$. For small initialization of the network’s weights, the alignment of $V(F_{T})$ with the eigenfunctions of $T_{k}$ depend on two factors: (i) alignment of $V(F_{0})$ with the eigenfunctions of $T_{k}$; (ii) singular values of $T_{k}$.
Under weak orthogonalization constraints, the network tends to learn features that are strongly aligned with eigenfunctions corresponding to large singular values. We refer to this property as the “selection” bias of gradient descent, wherein gradient descent selects certain eigenfunctions based on the corresponding singular values. This selection bias leads to redundancy among the learned feature space, thereby reducing the effective dimensionality of the network’s output space compared to its ambient dimensionality. We will leverage this finding to improve the parameter overhead of good feature learning using BarlowTwins and VICReg loss frameworks.
3.3 Takeaway 1: Low-dimensional projectors can yield good representations